Prompt: Simple and compounding interest – Given a 2-year simple interest account or 2-year compound interest...

  • Prompt: Simple and compounding interest – Given a 2-year simple interest account or 2-year compound interest account both paying 5% interest which would you rather receive? Discuss your preference and provide an example. Include in the discussion the difference between simple and compound interest.
  • Requirements: 250 words
Answer #1

Simple Interest as the name specifies only calculates the interest over the period and divides this interest among multiple periods. The interest will be on the initial principal.

In compound interest, interest is calculated over interest during a period ie interest will be calculated on the amount (principal +interest) after a particular period for the purpose of calculation for the next period.

As an example, consider a $ 1000 investment for 2 years at 5% interest(both simple and compound)

Simple Interest

Simple interest = P*N*R/100 = 1000*2*5 / 100 = $ 100

Amount after 2 years = $ 1,100

For Compound Interest

Amount in Compound Interest = 1000*(1+5%)^2 = $ 1,102.5

So Amount after compound interest for 2 years> Amount in SI

Thus compound interest is preferred over simple interest.